if 2x - 3y = -4 and 3x + 2y = 46 what are the values of x and y ? thankyou
Answer by uma(370)

Given 2x - 3y = - 4 --------------------(1)
3x + 2y = 46 -----------------------(2)
==> 4x - 6y = -8 --------------(1) x 2
9x + 6y = 138 ------------------(2) x 3
Adding the above equations we get,
13x = 130
==> 13x/13 = 130/13
==> x = 10
Plugging in this value in (2),
30 + 2y = 46
==> 2y = 46 - 30
==> 2y = 16
==> 2y/2 = 16/2
==> y = 8
Thus x = 10 and y = 8
Good Luck!!!